1f [Η3Ο+] = 1.0Χ 10-3,pOH = ? Ο 11 Ο Ο Ο
Answer #1

Given [H₂o+ ] = 1 x 10-3 m we know that PH+ Por = 14 - 0 and pr= -log [ht] or PH = -log [.zot Here a PH = -log Euz oth - – loany query please comment.

If satisfied please rate it thanks.